The graph of f is perpendicular to the line whose equation is 2 x - 3 y + 9 = 0 and has the same y-intercept as this line.
Answer by lwsshak3(11628)

The graph of f is perpendicular to the line whose equation is 2 x - 3 y + 9 = 0 and has the same y-intercept as this line.
**
2x-3y+9=0
3y=2x+9
y=2x/3+9/3
y=2x/3+3
This is an equation of a straight line of the standard form: y=mx+b, m=slope, b=y-inteercept
For given problem:
slope=2/3
y-intercept=3
..
For perpendicular line with same y-intercept
slope=negative reciprocal of slope of given equation=-3/2
Equation: f(x)=-3x/2+3